Yuk! I hate applications. I always get mixed up picking picking what goes where in the equations. I had one person show me how to solve it using just equations, another showed me how by using a table. I'm confused with both of them.
Here it is.
A truck rental agency charges a daily fee plus a milage fee. Mel was charged $85 for 2 days plus 100 miles, and Sarah was charged $165 for 3 days and 400 miles. What is the agency's daily fee and their milage fee?
a) Write the system of equations for this problem.
b) Find the solution.

Answer by Nate(3500)

x will be the cost per mile
y will be the cost per day
.
Mel was charged $85 for 2 days plus 100 miles.
85 bucks for the sum of the costs
2 days would cost 2y
100 miles would cost 100x
2y + 100x = 85
.
Sarah was charged $165 for 3 days and 400 miles.
165 bucks for the sum of the costs
400 miles would cost 400x
3 days would cost 3y
3y + 400x = 165
.
a) Write the system of equations for this problem.
2y + 100x = 85
3y + 400x = 165
.
b) Find the solution.
-4(2y + 100x = 85)
3y + 400x = 165
.
-8y - 400x = -340
3y + 400x = 165
.
-5y = -175
y = 35
.
Pick one of the initial equations:
2y + 100x = 85
2(35) + 100x = 85
70 + 100x = 85
100x = 15
x = 0.15
.
$35 per day and $0.15 per mile